OpenStudy (anonymous):

y = negative three-halvesx + 4

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

identify the slope and y intercept

OpenStudy (anonymous):

In y=mx+b, m=slope, b=y intercept

OpenStudy (anonymous):

That form of the equation is called "slope-intercept form" for that reason.

OpenStudy (anonymous):

ok now what is the answer

OpenStudy (anonymous):

That's a good question.

OpenStudy (paxpolaris):

in \(\large y=mx+b\), m is slope, and b is y-intercept: So, in\[\Large y= -\frac32x+4\].........................................................

OpenStudy (paxpolaris):

give an answer... we'll tell you if you ar on the right track...

OpenStudy (anonymous):

Y-intercept=4 Slope= 3/2

OpenStudy (paxpolaris):

y-intercept is 4 ... YES but, slope is \(-3/2\).... remember to keep the negative

OpenStudy (anonymous):

k thanks